# Segregation in Nuclear Stellar Clusters: Rates and Mass Distributions of TDEs, QPEs, Plunges, and EMRIs

### v1

@@ 2.2 BHs mass function
We assume that BHs at the mass range $10\lesssim m_{\bullet}\lesssim 30$ follow a power law profile
@@
As our fiducial values, we adopt $\gamma_{\bullet}=3$, a BH number fraction $f^{\rm IMF}_{\bullet}=10^{-3}\approx N_{\bullet}(10)/N_{\star}(1)$, as expected from the stellar IMF, and, consequently, $\xi_{\bullet}=0.1$. This distribution was chosen as a simple example in which the $30{M_{\odot}}$ BHs dominate the scattering. Our qualitative analysis, therefore, applies more broadly for BH distributions that satisfy this condition (as further discussed in §2.4).
We assume that the black hole mass distribution declines significantly outside the range $10\lesssim m_{\bullet}\lesssim 30$ and has little influence on our results. At the lower mass end, we adopt a minimal BH mass of $m_{\bullet}=5$ (consistent with the “lower mass gap”, see Abbott et al., 2023). This distribution is schematically presented in Fig. (2).
Figure 1: The enclosed mass of stars (solid colored lines) and BHs (dashed and dashed-dotted lines) as a function of semimajor axis $r$. The purple dashed vertical line marks $R_{\bullet}$ (Eq. 11), below which $30{M_{\odot}}$ BHs dominate the scattering and follow the BW profile. In this region, lighter stars and BHs follow shallower power-law profiles. At larger distances ($R_{\bullet}\lesssim r\lesssim R_{h}$), massive objects become exponentially rare as they segregate inward (as derived by Linial & Sari, 2022). The BH distribution flattens around $R_{h}$ due to the influx of BHs from greater distances. The red arrow indicates the observational upper limit for the extended mass enclosed within the orbit of S2 (Gravity Collaboration et al., 2024). The distributions are truncated once a single object remains.
@@ 2.3 Stellar Segregation
Assuming that the nuclear stellar cluster is dynamically relaxed, the stellar steady-state distribution satisfies a zero particle flux (Bahcall & Wolf, 1976, 1977; Linial & Sari, 2022), which corresponds to a constant, non-vanishing energy flux (Binney & Tremaine, 1987; Fragione & Sari, 2018; Rom et al., 2023).
Linial & Sari (2022) derived the zero-flux, steady-state solution for a power-law mass function and observed that weighting the segregated distribution by a factor of $m^{3/2}$ reproduces the single-mass BW profile. Here, we offer a simple explanation for this result, which allow us to apply it for the more realistic PMF, as given by Eq. (4).
@@
where $E(r,m)\propto m/r$ is the orbital energy, and $\tau_{\rm rlx}$ is the relaxation timescale. Furthermore, $\tau_{\rm rlx}$ is given by (Binney & Tremaine, 1987; Merritt et al., 2011; Sari & Fragione, 2019)
@@
Combining Eqs. (7) and (8) we find the energy flux scales as $\mathcal{F}_{E}\propto m^{3}N^{2}r^{-5/2}$. Maintaining a constant energy flux, with a given mass spectrum (e.g., Eqs. 4 and 5), determines the density profile as well as the mass that dominates the scattering at each radius. In this interpretation, the effective cross-section of $m^{3/2}$, identified by Linial & Sari (2022), arises from the combination of the scattering cross-section, $\sigma\propto m^{2}$, and the orbital energy, $E\propto m$.
The constant energy flux solution naturally leads to mass-segregated distribution. Massive stars (with $m\gtrsim 1$) sink inward up to a characteristic distance $r_{d}(m)$ where their self-scattering dominate the energy flux. Using the mentioned scaling, together with the constant energy flux requirement and the stellar PMF, gives
@@
The spatial distribution of different stars and BHs is demonstrated in Fig. (1), where we present the enclosed mass as a function of the semimajor axis. Evidently, low-mass stars ($m\lesssim 1$) dominate the mass budget throughout most of the cluster. The contribution from more massive objects to the total enclosed mass becomes comparable to that of the low-mass stars around $r\sim 10^{-4}R_{h}$, where only a few of these massive objects reside.
Recent observations of the Milky Way’s center (Gravity Collaboration et al., 2024) impose an upper limit of about $1200{M_{\odot}}$ for the extended mass within the orbit of S2 (depicted by the red arrow in Fig. 1). Our segregated profile predicts an extended mass value similar to this upper limit, composed mostly of solar-mass stars. Notably, around S2’s semimajor axis, the density of $10{M_{\odot}}$ BHs is a factor of $\approx 30$ lower than that of the stars. Furthermore, without mass segregation, a BW profile of stars would predict a stellar density higher by a factor of $\approx 5$, in contrast with the observations.
2.4 BHs segregation
The assumption of a relaxed nuclear stellar cluster implies that massive objects can migrate into the sphere of influence during the galaxy lifetime. Beyond the radius of influence, the stellar distribution can be approximated as an isothermal sphere, where $N(r)\propto r$ and the relaxation timescale for objects of mass $m$ scales as $\tau_{\rm rlx}(r>R_{h},m)\propto r^{2}/m$ (Binney & Tremaine, 1987).
Therefore, more massive objects can segregate inwards from greater distances, scaling as $r\propto m^{1/2}$. This process increases the mass distribution of stellar mass black holes within the sphere of influence by a factor of $m_{\bullet}^{1/2}$, resulting in an effective power-law index
@@ 2.4 BHs segregation
Thus, the BH number fraction inside the sphere of influence, $f^{\rm cusp}_{\bullet}$, is expected to be enhanced by $f^{\rm cusp}_{\bullet}/f^{\rm IMF}_{\bullet}\approx 3$ relative to its expected value from the (non-segregated) stellar IMF. This enhancement impacts the expected number of EMRIs that would be observed by LISA (Babak et al., 2017; Bonetti & Sesana, 2020; Pozzoli et al., 2023; Rom et al., 2024). Additionally, this incoming flux of BHs flattens their spatial distribution near the radius of influence, as schematically illustrated in Fig. (1).
@@
Within the sphere of influence, the BHs sink, due to dynamical friction from the stars, to a characteristic distance $R_{\bullet}$, where BHs with masses $\widetilde{m}_{\bullet}\approx 30$ become the dominant scatterers. This distance is determined by comparing the energy flux of BHs (analogous to Eq. 7) to the constant energy flux set by one solar-mass stars at the radius of influence. This yields
@@
As presented in Fig. (1), from $R_{\bullet}$ inward, the $30{M_{\odot}}$ BHs follow the single-mass BW profile, while lighter stars and BHs settle into shallower profiles (according to Eq. 1). Thus, massive stars with $m\gtrsim 3$ will accumulate around $R_{\bullet}$, where they will be scattered by the BHs, rather than efficiently segregating to $r_{d}(m)\lesssim R_{\bullet}$. The critical stellar mass is given by $\widetilde{m}=\xi_{\bullet}^{2/\left(5-2\gamma\right)}\approx 3$, satisfying $r_{d}(\widetilde{m})\approx R_{\bullet}$. Therefore, we conclude that in the segregated distribution, stars of mass $m$ are concentrated around a characteristic distance, $\widetilde{r}_{d}(m)$, given by
@@
where $r_{d}(m)$ and $R_{\bullet}$ are given in Eqs. (9) and (11), respectively.
@@
Our results are weakly sensitive to the poorly known stellar-mass BH distribution as long as heavy $\approx 30{M_{\odot}}$ BHs are abundant enough to dominate the scattering, namely, $N_{\bullet}(30)/N_{\bullet}(10)\gtrsim 3^{-3/2}$. The observations of the LIGO-Virgo-KAGRA (LVK) collaboration (Abbott et al., 2023) suggest that this is indeed the case, if the binary merger population is a good representative of the overall mass distribution of stellar-mass BHs.
We consider a BH mass function with $\gamma_{\bullet}=3$, leading to $\widetilde{\gamma}_{\bullet}=5/2$, which represents the steepest power-law profile where $30{M_{\odot}}$ BHs dominate the scattering. We note that for steeper profiles, $\gamma_{\bullet}>3$, with the same BH number fraction, the $10{M_{\odot}}$ BHs become the dominant scatterers at $R_{\bullet}$, and the impact of more massive stellar-mass BHs diminishes, as they are rarer and more segregated. Namely, the more massive stellar-mass BHs may dominate the scattering only at smaller distances, $r<R_{\bullet}$, where fewer objects reside.
@@ 3 Transients mass-dependent rates
Based on the steady-state distributions of stars and BHs, the transients formation rates can be estimated using the two-body scattering induced flux (Wang & Merritt, 2004; Hopman & Alexander, 2005; Stone & Metzger, 2016). Here we generalize this concept by taking into account the mass segregation as a function of radius, and arrive at
@@
We estimate the rate of each transient type and object mass by evaluating Eq. (13) at the radius where it is maximal, as described below. Additionally, we define the enhancement factor $\mathcal{Q}$ as the ratio between the rate of a specific transient, involving stars or BHs, and their abundance in the total population

### v2

@@ 2.2 BH mass function
We assume that BHs at the mass range $10\lesssim m_{\bullet}\lesssim 30$ follow a power-law profile
@@
As our fiducial values, we adopt $\gamma_{\bullet}=3$, a BH number fraction $f^{\rm IMF}_{\bullet}=10^{-3}\approx N_{\bullet}(10)/N_{\star}(1)$, as expected from the stellar IMF, and, consequently, $\xi_{\bullet}=0.1$. This distribution was chosen as a simple example in which the $30{M_{\odot}}$ BHs dominate the scattering. However, our qualitative results do not depend on the exact details of the BH mass function, as discussed in Section 2.4.
We assume that the BH mass distribution declines significantly outside the range $10\lesssim m_{\bullet}\lesssim 30$ and has little influence on our results. At the lower-mass end, we adopt a minimal BH mass of $m_{\bullet}=5$ (consistent with the “lower-mass gap”, see Abbott et al., 2023). This distribution is schematically presented in Figure (2).
Figure 1: The enclosed mass of stars (solid colored lines) and BHs (dashed and dashed-dotted lines) as a function of semimajor axis $r$. The purple dashed vertical line marks $R_{\bullet}$ (Equation 11), below which $30{M_{\odot}}$ BHs dominate the scattering and follow the BW cusp. In this region, lighter stars and BHs follow shallower power-law profiles. At larger distances ($R_{\bullet}\lesssim r\lesssim R_{h}$), massive objects become exponentially rare as they segregate inward (as derived by Linial & Sari, 2022). The BH distribution flattens around $R_{h}$ due to the influx of BHs from greater distances. The red arrow indicates the observational upper limit for the extended mass enclosed within the orbit of S2 (Gravity Collaboration et al., 2024). The distributions are truncated once a single object remains.
@@ 2.3 Stellar Segregation
Assuming that the nuclear stellar cluster is dynamically relaxed, the stellar steady-state distribution satisfies a zero particle flux (Bahcall & Wolf, 1976, 1977; Linial & Sari, 2022), which corresponds to a constant, nonvanishing energy flux (Binney & Tremaine, 1987; Fragione & Sari, 2018; Rom et al., 2023).
Linial & Sari (2022) derived the zero-flux, steady-state solution for a power-law mass function and observed that weighting the segregated distribution by a factor of $m^{3/2}$ reproduces the single-mass BW cusp. Here, we offer a simple explanation for this result, which allow us to apply it for the more realistic PMF, as given by Equation (4).
@@
where $E(r,m)\propto m/r$ is the orbital energy, and $\tau_{\rm rlx}$ is the relaxation timescale. Furthermore, $\tau_{\rm rlx}$ is given by (Binney & Tremaine, 1987; Merritt et al., 2011)
@@
Combining Equations (7) and (8) we find the energy flux scales as $\mathcal{F}_{E}\propto m^{3}N^{2}r^{-5/2}$. Maintaining a constant energy flux, with a given mass spectrum (e.g., Equations 4 and 5), determines the density profile as well as the mass that dominates the scattering at each radius. In this interpretation, the effective cross section of $m^{3/2}$, identified by Linial & Sari (2022), arises from the combination of the scattering cross-section, $\sigma\propto m^{2}$, and the orbital energy, $E\propto m$.
The constant energy flux solution naturally leads to mass-segregated distribution. Massive stars (with $m\gtrsim 1$) sink inward up to a characteristic distance $r_{d}(m)$ where their self-scattering dominates the energy flux. Using the mentioned scaling, together with the constant energy flux requirement and the stellar PMF, gives
@@
The spatial distribution of different stars and BHs is demonstrated in Figure (1), where we present the enclosed mass as a function of the semimajor axis. Evidently, low-mass stars ($m\lesssim 1$) dominate the mass budget throughout most of the cluster. The contribution from more massive objects to the total enclosed mass becomes comparable to that of the low-mass stars around $r\sim 10^{-4}R_{h}$, where only a few of these massive objects reside.
Recent observations of the Milky Way’s center place an $1\sigma$ upper limit of approximately $1200{M_{\odot}}$ on the extended mass enclosed within the orbit of S2 (depicted by the red arrow in Figure 1; Gravity Collaboration et al., 2024). Our segregated profile predicts an extended mass of about $\sim 2000{M_{\odot}}$, slightly exceeding this observational constraint. However, we expect stellar collisions to deplete the number of stars in the inner regions of the cluster, thereby lowering the extended mass within the S2 orbit. A detailed account of the impact of collisions is deferred to a subsequent paper.
2.4 BH Segregation
The assumption of a relaxed nuclear stellar cluster implies that massive objects can migrate into the sphere of influence during the galaxy’s lifetime. Beyond the radius of influence, the stellar distribution can be approximated as an isothermal sphere, where $N(r)\propto r$ and the relaxation timescale for objects of mass $m$ scales as $\tau_{\rm rlx}(r>R_{h},m)\propto r^{2}/m$ (Binney & Tremaine, 1987).
Therefore, more massive objects can segregate inwards from greater distances, scaling as $r\propto m^{1/2}$. This process increases the mass distribution of stellar-mass BHs within the sphere of influence by a factor of $m_{\bullet}^{1/2}$, resulting in an effective power-law index
@@ 2.4 BH Segregation
Thus, the BH number fraction inside the sphere of influence, $f^{\rm cusp}_{\bullet}$, is expected to be enhanced by $f^{\rm cusp}_{\bullet}/f^{\rm IMF}_{\bullet}\approx 3$ relative to its expected value from the (nonsegregated) stellar IMF. This enhancement impacts the expected number of EMRIs that would be observed by LISA (Babak et al., 2017; Bonetti & Sesana, 2020; Pozzoli et al., 2023; Rom et al., 2024). Additionally, this incoming flux of BHs flattens their spatial distribution near the radius of influence, as schematically illustrated in Figure (1).
@@
Within the sphere of influence, the BHs sink, due to dynamical friction from the stars, to a characteristic distance $R_{\bullet}$, where BHs with masses $\widetilde{m}_{\bullet}\approx 30$ become the dominant scatterers. This distance is determined by comparing the energy flux of BHs (analogous to Equation 7) to the constant energy flux set by one solar-mass stars at the radius of influence. This yields
@@
As presented in Figure (1), from $R_{\bullet}$ inward, the $30{M_{\odot}}$ BHs follow the single-mass BW cusp, while lighter stars and BHs settle into shallower profiles (according to Equation 3). Thus, massive stars with $m\gtrsim 3$ will accumulate around $R_{\bullet}$, where they will be scattered by the BHs, rather than efficiently segregating to $r_{d}(m)\lesssim R_{\bullet}$. The critical stellar mass is given by $\widetilde{m}=\xi_{\bullet}^{2/\left(5-2\gamma\right)}\approx 3$, satisfying $r_{d}(\widetilde{m})\approx R_{\bullet}$. Therefore, we conclude that in the segregated distribution, stars of mass $m$ are concentrated around a characteristic distance, $\widetilde{r}_{d}(m)$, given by
@@
where $r_{d}(m)$ and $R_{\bullet}$ are given in Equations (9) and (11), respectively.
@@
Our results are weakly sensitive to the shape of the poorly constrained BH mass function. In our fiducial model, we assume that the heavy $\approx 30{M_{\odot}}$ BHs are sufficiently abundant to dominate the scattering, namely, $N_{\bullet}(30)/N_{\bullet}(10)\gtrsim 3^{-3/2}$. The observations of the LIGO-Virgo-KAGRA (LVK) collaboration (Abbott et al., 2023) suggest that this is indeed the case, if the binary merger population is a good representative of the overall mass distribution of stellar-mass BHs. Specifically, we adopt a power-law BH mass function with $\gamma_{\bullet}=3$ (which leads to an effective slope $\widetilde{\gamma}_{\bullet}=5/2$ due to sinking of BHs from outside the sphere of influence, see Equations 6 and 10). This is the the steepest power-law profile where $30{M_{\odot}}$ BHs dominate the scattering. Therefore, a shallower profile would result in limited changes to our results.
However, if the more massive stellar-mass BHs are less abundant, e.g., a steeper mass function with $\gamma_{\bullet}>3$, then the $10{M_{\odot}}$ BHs become the dominant scatterers at $R_{\bullet}$, assuming the total BH number fraction remains fixed. In this case, the scattering timescale at $r\lesssim R_{\bullet}$ increases by a factor of $3^{1/2}$, compared to our fiducial model (see Equation 15). The impact of heavier stellar-mass BHs then becomes subdominant, as they concentrate at smaller radii, $r<R_{\bullet}$, where fewer objects reside and their contribution to the overall dynamics is diminished.
2.5 Comparison with numerical studies
Our result for the stellar and BH distributions, derived from the condition of constant energy flux (as presented in Figure 1), indicates that stellar-mass BHs sink and become the dominant scatterers within $R_{\bullet}\sim 0.1R_{h}$ (Equation 11). In this inner region ($r\lesssim R_{\bullet}$), the BHs form a BW cusp, with a power-law slope $\eta_{\bullet}=-(d\log n/d\log r)=1.75$, while lighter objects, with mass $m_{\ell}$, develop shallower profiles, $\eta_{\ell}=1.5+m_{\ell}/(4m_{\bullet})\approx 1.5$, as given by Equation (3).
Dynamically, we expect somewhat flatter profiles than noted above, especially in the inner regions of the cluster, if there is a sink term, such as stellar collisions or loss-cone depletion (Cohn & Kulsrud, 1978; Bar-Or & Alexander, 2016), which is not included in our model. Observationally, such flattened profiles, with slopes $\eta=1.1-1.4$, have been reported for the stellar population in our own Galactic center (Gallego-Cano et al., 2018; Schödel, R. et al., 2018).
Several numerical studies have found that BHs develop a BW cusp, while lighter objects exhibit shallower profiles, in the inner regions of the nuclear stellar cluster ($r\lesssim 0.1R_{h}$), in agreement with our prediction. Table 1 summarizes the power-law slopes inferred by the various numerical studies discussed below.
Freitag et al. (2006) used a spherical stellar dynamical MC code that includes weak and strong scattering, stellar evolution, collisions, and an approximate treatment of loss-cone depletion. Their results show the expected cusp slopes for both the BHs and the stars (see their Figs. 8 and 10 for two- and multimass models, respectively). Broggi et al. (2022) obtained similar slopes using a FP approach for a two-mass system of $1{M_{\odot}}$ stars and $10{M_{\odot}}$ BHs (which we refer to as the fiducial two-mass system), including scattering and loss-cone depletion (see their Figure 4). Panamarev et al. (2019) conducted direct N body simulations with one million particles, accounting for stellar evolution and binary stars, and found consistent slopes for BHs, WDs, and stars (see their Figure 3). Baumgardt et al. (2018) also employed a direct N body simulation (with $5\times 10^{4}$ particles), including scattering, stellar evolution, and star formation. While their results show that BHs follow the steepest profile, the inferred slope ($\eta_{\bullet}=1.55$) is shallower than values reported in other studies. This discrepancy may be due to the fact that the reported slope reflects an average over both BHs and lighter neutron stars (NSs).
Balberg (2024) further confirmed these trends through MC simulations of the fiducial two-mass system, incorporating weak scattering, GW emission, loss-cone depletion, and stellar collisions (see their Figs. 3 and 4). Their results support two additional predictions of our model (and that of Rom et al., 2024, for the two-mass case). First, stars near $R_{h}$ follow a BW cusp, whereas BHs develop a steeper profile, although their reported slope of $\eta_{\bullet}\approx 2$ is shallower than our analytical prediction (see discussion below). Second, they confirm that the radial extent of the BH cusp depends on the BH number fraction $f_{\bullet}$, in excellent agreement with our analytical estimate (see Table 1, Equation (11) here and Equation (10) in Rom et al., 2024).
Other numerical studies find steeper cusps for the BHs, $\eta_{\bullet}\approx 2$. For instance, Hopman & Alexander (2006) numerically solved the FP equation for a multimass population composed of stars, WDs, NSs, and BHs (see their Figure 1). O’Leary et al. (2009) extend this approach to models with power-law BH mass function (see their Figure 1). They found that BHs dominate within $0.1R_{h}$, where the lighter components follow shallower profiles consistent with Equation (3), though with $p_{\max}\approx 0.5$ rather than $p_{\max}\approx 0.25$ (as in Bahcall & Wolf, 1977).
Preto & Amaro-Seoane (2010) carried out N body simulation (with $\sim 10^{5}$ particles) and FP calculation for the fiducial two-mass system, approximately reproducing either the Bahcall & Wolf (1977) solution or the Alexander & Hopman (2009) solution depending on the BH number fraction (see their Figure 2). Aharon & Perets (2016) applied the FP method to models with different star formation histories, including stars, WDs, NSs, $10{M_{\odot}}$ and/or $30{M_{\odot}}$ BHs. They found that BHs consistently develop the steepest profile, with the exact slope depending on the assumed star formation scenario (see their Figs. 1 and 2) [^3] An exception is a model including in situ star formation beyond the sphere of influence, in which all species, including BHs, exhibit flatter slopes of $\eta\approx 1$.. Finally, Zhang & Amaro Seoane (2025) explored a suite of two-mass models using an MC method, incorporating two-body scattering, the evolving stellar potential, and SMBH growth. Their results are consistent with the general trend of steep BH cusps and shallower profiles for lighter species (see their Figure 9).
The discrepancies in the reported slopes, as presented in Table 1, may arise from limitations in numerical resolution or small number statistics, as the innermost regions of the cluster are sparsely populated. Additionally, in the outer region of the cluster, $R_{\bullet}\lesssim r\lesssim R_{h}$, our model predicts a very steep profile for the massive objects. In the multimass case, they become exponentially rare (Linial & Sari, 2022), while in a two-mass model their distribution follows a power law with $\eta_{\bullet}=4$ for $10M_{\odot}$ (Rom et al., 2024), and even steeper slopes for more massive BHs, following Equation 3. These predictions are steeper than those of Alexander & Hopman (2009) in their “strong segregation” regime, as we argue for a constant energy flux while they assume a constant particle flux.
To our knowledge, such steep stellar-mass BH profiles near the radius of influence have not been observed in numerical simulations. This may be due to the fact that these extreme slopes occur only within a narrow radial range (as illustrated in Figure 1), making them challenging to resolve numerically. Nevertheless, the presence of steep slopes in the cluster outskirts could bias the inferred average slope if the BH distribution is fitted with a single power law over the entire sphere of influence.
Table 1: Comparison of BH and stellar distributions in nuclear stellar clusters from numerical studies
Reference $\eta_{\bullet}$aaSlope of the distribution, $\eta_{\bullet}=-(d\log n_{\bullet}/d\log r)$, for the dominant stellar-mass BHs. $\eta_{\ell}$bbSlope of the distributions of the lighter populations, which may include stars, WDs, and NSs, depending on the specific model. Radial ExtentccRadial extent over which the reported power-law slopes are estimated.
This work 1.75 $\approx$ 1.5 $r\lesssim 0.16R_{h}$
Freitag et al. (2006) 1.71.81.7-1.8 1.31.41.3-1.4 $r\lesssim 0.15R_{h}$
Hopman & Alexander (2006) $\approx 2$ 1.41.51.4-1.5 $r\lesssim R_{h}$
O’Leary et al. (2009) $\approx 2$ $\approx 1.5$ $r\lesssim 0.1R_{h}$
Preto & Amaro-Seoane (2010) 2.1(1.8)2.1\ (1.8)ddValue depends on the BH number fraction, $f_{\bullet}\approx 2.5\times 10^{-3}$ ($f_{\bullet}\approx 0.5$). $1.5$ $r\lesssim R_{h}\ (0.1R_{h})$ddValue depends on the BH number fraction, $f_{\bullet}\approx 2.5\times 10^{-3}$ ($f_{\bullet}\approx 0.5$).
Aharon & Perets (2016) 1.92.31.9-2.3 1.31.51.3-1.5 $r\lesssim 10^{-2}R_{h}$
Baumgardt et al. (2018) $1.55$ 11.21-1.2 $r\lesssim R_{h}$
Panamarev et al. (2019) $1.72\pm 0.04$ 0.91.20.9-1.2 $r\lesssim 0.5R_{h}$
Broggi et al. (2022) $1.7$ $1.3$ $r\sim 10^{-3}R_{h}$
Balberg (2024) $1.75$ $1.475$ $r\lesssim 0.05R_{h}\ (0.2R_{h})$eeValue depends on the BH number fraction, $f_{\bullet}\approx 10^{-3}$ ($f_{\bullet}\approx 10^{-2}$).
Zhang & Amaro Seoane (2025) 1.752.21.75-2.2 1.11.41.1-1.4 $0.01R_{h}\lesssim r\lesssim R_{h}$
@@ 3 Transients mass-dependent rates
Based on the steady-state distributions of stars and BHs, the transient formation rates can be estimated using the two-body scattering induced flux (Wang & Merritt, 2004; Hopman & Alexander, 2005; Stone & Metzger, 2016). Here, we generalize this concept by taking into account the mass segregation as a function of radius, and arrive at
@@
We estimate the rate of each transient type and object mass by evaluating Equation (13) at the radius where it is maximal, as described below. Additionally, we define the enhancement factor $\mathcal{Q}$ as the ratio between the rate of a specific transient, involving stars or BHs, and their abundance in the total population